OpenStudy (anonymous):

Need help with the last 2. Can't seem to find the answer choices they gave. Keep coming up with a weird fraction log answer. I'm trying to assist my child with this. How do I input this in her Ti83 calculator? Thanks

OpenStudy (anonymous):

Can you upload a screen shot of the actual problem?

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

There must have been a problem with the upload since I'm not seeing any attached files

OpenStudy (anonymous):

Trying again

OpenStudy (sloppycanada):

I assume the numbers on the side is the answer that she's chosen or are they our options?

OpenStudy (anonymous):

Yes those are the answer choices.

OpenStudy (campbell_st):

well you need to know a few log laws \[2\log_{10}(6) ~`can~be ~written ~~\log_{10}(6^2)\] same thing for the other term \[\frac{1}{3} \log_{10} 27 = \log_{10} \sqrt[3]{27}\] so if you simplify the terms you get \[\log_{10}(36) - \log_{10}(3)\]

OpenStudy (anonymous):

The first 3. She got the answers 4, 3 & 2

OpenStudy (campbell_st):

with logs... subtracting the terms really means you are dividing the logs \[\log_{10}(\frac{36}{3}) = \log_{10}(x)\] so to find x, simplify 36/3

OpenStudy (campbell_st):

is the last question also a problem... ?

OpenStudy (campbell_st):

so what you are being asked to do, is get a single log on the left side of the equation so that you can then solve for x

OpenStudy (anonymous):

Thanks. I'll go over this with her tomorrow & see if it makes better sense to her.

OpenStudy (campbell_st):

ok as for the last question the log law you need is \[\log_{a}(\frac{m}{n}) = \log_{a}(m) - \log_{a}(n)\] so this means that the left side of the equation can be written as \[\log_{7}(24) - \log_{7}(x + 5) = \log_{7}(\frac{24}{x + 5})\] so the problem becomes \[\log_{7}(\frac{24}{x + 5}) = \log_{7}(8)\] since they are the same base you can then equate \[\frac{24}{x + 5} = 8\] which becomes \[24 = 8(x + 5)\] now you can solve for x divide both sides of the equation by 8 then subtract 5. hope it helps

OpenStudy (anonymous):

Very much! So glad you broke it down. Makes more sense now